# The position and resolvability of blended point sources

### v1

@@ 1 Introduction
This applies to pairs of sources that are truly close in physical space (e.g. perhaps a gravitationally bound binary system) or that are at very different distances from the observer and each other but appear close in projection. It is also eminently possible for more than two sources to be blended - for example whole clusters of stars appear as a bright point source in extragalactic observations (Harris91). In fact we may expect many blended sources to be contributing to an observed image (for example planets and asteroids around a star, or many more distant background stars) though if these systems contribute a negligible fraction of the observed light, their influence is small.
@@
Binary stars (and higher multiples) are ubiquitous in most astrophysical environments. Only very wide systems, or those very close to the observer, are resolvable - however multiplicity affects many facets of their behaviour and evolution. To some they are objects of interest in their own right, to others a nuisance to be removed or compensated for when making observations. Only a subset can be identified, and only a subset of those can have their properties measured. The number of known multiples is rapidly increasing, both by resolving wide systems (for example Tokovinin24) or, for unresolved systems, by identifying variable or excess motion (spectroscopically or astrometrically, see El-Badry24 for a recent review) or light (e.g. Prsa05 or Wallace24).
The Gaia survey has been crucial to the improvement in our capacity to detect stellar multiples, as well as greatly expanding our knowledge of the position and motion of stars in the Milky Way (Prusti16; Vallenari23). It is, at least in part, an astrometric survey - scanning repeatedly over the sky and making many measurements of the positions of sources, from which their motion can be inferred (Lindegren21). In Penoyre20 we mapped out how the orbits of unresolved binaries contribute to this motion, due to the offset between their photocentre (or center-of-light) and their center-of-mass. In appendix A of that work, we derived the simple first order photocentre offset distance, without including an important nuance to Gaia’s functionality: that its mirror is rectangular, measuring 1.45 m by 0.5 m (Prusti16), and hence its subsequent PSF is extended along the direction corresponding to the short axis of the mirror.
@@ 2 Brightness profiles
Throughout this work we will work with Gaussian point-spread functions (PSF). These are vastly simplified from real diffraction patterns. Idealised forms of elongated PSFs include convolutions of two $\mathrm{sinc}^{2}$ functions for a rectangular aperture, or an elongated form of the well known Airy disk for an elliptical aperture (Airy35; Raman19).
@@
True PSFs will be further altered by any non-ideal aberrations to the mirror, optical path, or detector. For example Fabricius16 show a typical Gaia PSF (figure 7) which as expected resembles a smoothed $\mathrm{sinc}^{2}$ function, but displays significant asymmetry and strong diagonal spikes. Real observations are then further altered by pixelation, losing more information as the incident flux is binned over a finite number of pixels. Very bright sources can oversaturate these pixels, spilling charge and thus apparent brightness into neighbouring pixels, and further deforming the PSF.
@@ 2.1 Imaging close sources
For derivations in this paper we’ll work entirely in units of $\alpha$ and $\beta$ (and hence we shall not need numerical values specific to a given instrument). However, as Gaia is the most currently relevant point of comparison we note that its full-width-half-maximum is approximately 0.1 arcseconds (Fabricius16) in the along scan direction, and hence $\alpha\sim 0.04$ arcseconds. Given that the ratio of side lengths of the mirror is 1.45:0.5 we expect $\beta\sim 3\alpha$.
@@
Real observations may use a more complex PSF fitting technique (e.g. Rowell21; Harrison23), but the above assumptions provide a reasonable approximation that is well defined mathematically.
@@ 2.2 Maxima and minimum
Equations 14 and 15 are the equivalent of equation 64 in Lindegren22. However, their calculation only includes the along-scan PSF width, $\alpha$ ($u$ in their notation), implicitly assuming $\beta\gg\alpha$ (and hence $\gamma\rightarrow\alpha/|\mathrm{c}_{\phi}|$). This significantly overestimates the effective width as $|\mathrm{c}_{\phi}|\rightarrow 0$.
@@ 2.2.1 Approximate forms
where $x_{0}$ is the much simplified zeroth order equation for the photocentre reported in Penoyre20.
@@
$x_{m}=x_{0}\left(1+\frac{x_{0}^{2}}{l\gamma^{2}}\right)+\mathcal{O}\left(\frac{r^{4}}{\gamma^{4}}\right).$ (18)
@@ 3.1 Numerical solutions
We seek the solution(s), $x_{m}$, of equation 13. This can be expressed as root(s) of $g(\chi_{m})=0$ which can be reliably found by any root-solver which operates over a range $a<\chi<b$ such that $g(a)$ and $g(b)$ have opposite sign and there is a single root in this range (for example the van Wijngaarden-Deker-Brent method, Brent02 - implemented as brentq in scipy.optimize).
@@
The numerically derived position of each root is shown for a variety of $l$ and $\rho$ in figure 4. For small separations ($\rho<1$) the primary maximum is located close to the first order solution as found in Penoyre20: $x_{m}\sim x_{0}$. As $\rho$ increases the relative offset decreases, fastest for smaller $l$ - thus the first order solution is a good approximation for small $\rho$ and an overestimate for $\rho\gtrsim$ 1. In appendix A we derive the conditions that maximise $x_{1}$, as is also shown in the figure. At low $l$ the photocenter offset is maximised at $\rho=1$, and as $l$ increases this goes to $\rho=2$.
@@ 4 Repeated observations of a fixed pair of sources
We also show the across-scan position that one would find if $\beta\rightarrow\infty$ - as is implicitly assumed by the calculation in Lindegren22. This would lead to greater predicted offsets, especially at large separations, as any source observed at an angle such that $|\mathrm{c}_{\phi}|\lesssim\alpha/r$ would appear to be completely blended.
@@ 4.1 Binary offset and error
The precision to which the position of a source can be measured depends on the width of the PSF and how well sampled it is. For $N$ samples (i.e. measured photons) of a distribution with PSF width $\sigma$ we might expect a positional uncertainty of $\sim N^{-\frac{1}{2}}\sigma$. Other observational constraints, such as saturation of CCD pixels, can limit the possible precision. Lindegren21 (specifically appendix A) detail the current along-scan astrometric precision of Gaia, showing that it is photon noise dominated for systems with apparent magnitude $m_{G}\gtrsim 14$, and for brighter sources is approximately constant, with $\sigma_{al}\sim 0.2$ milliarcseconds.
@@ 4.2 Observational expectations
Assuming a single source (as the main Gaia source catalogue does) this motion is a combination of proper motion and parallax. If the scans are frequent and randomly oriented we would not expect any consistent behaviour that could mimic astrometric motion, thus we will only consider the position offset in detail in this work. However if there is some aliasing or coherent time evolution of scan-angles and times the varying positional offset may mimic motion, as seen in Holl23.
If there is excess unmodelled astrometric motion, for example from binarity, there will be a poor astrometric fit. Due to the ubiquity of binaries and higher multiples (accounting for around half of all systems Raghavan10) persistently poor fits can be used to identify probable multiple systems (see e.g. Penoyre20; Belokurov20; Penoyre22b).
@@
In Gaia DR3 a system can be distinguished as a significantly bad fit (whether due to internal motion or other contaminants) if the excess $RUWE$ is $\gtrsim 0.25$ and thus (Penoyre22; Castro-Ginard24).
@@ 4.2.1 Observable offset and error
We see the largest deviations from good astrometric fits, $\sqrt{1+D}-1$, at high light ratios and in the partially resolved region. The excess noise can be very large, well above the threshold of 0.25 appropriate for current Gaia data. Even excluding the semi-resolved region excesses of up to 50 are possible. This is a relatively hard predicted upper limit, and agrees well with the maximum RUWE seen in sources separated by around 1 arcseond (Belokurov20, figure 5).
@@ 4.2.2 Observational complications
Systems in the offset and error regime are the most amenable to detection with other methods, more observations or in later surveys. The large light ratios make them easier to discern from single sources, either from elongated photometry, or two discernable profiles in their spectra. Gaia includes flags based on the Image Parameter Determination (IPD) pipeline that record deviations from expected point-source brightness profiles which can be used to flag some sources in this regime (see for example appendix B of Penoyre22b). The pipeline for Gaia DR4 will identify and fit secondary peaks where possible, though the regime of interest here is at the edge of, or below, the threshold of resolvability and thus this will mitigate the effects mostly of the partially resolved systems.
@@ 5 Conclusions
We start, in section 2, with an elongated Gaussian PSF, as is appropriate for surveys such as the Gaia mission, with a rectangular focussing mirror. We show that this more general case can be reduced to a 1D problem, and thus converges with the circular PSF case when normalised by a scan-angle dependant effective PSF width, $\gamma(\phi)$ (equation 16). Thus any general case can be expressed in terms of $\rho=\frac{r}{\gamma}$, where $r$ is the on-sky separation of the sources. This introduces a further dependence on the relative orientation of the sources and our instrument, and shows that when this orientation changes over multiple measurements the variation in observed photocentre can provide extra information about the relative separation and light ratio of the sources. Such variations have indeed been observed in Gaia already, as detailed in Holl23.
@@
When two blended sources maintain a fixed orientation analytic expressions for their position offset and induced error can be derived, as we do in section 4. This is a particular case of interest for long period (and hence effectively stationary) binaries observed at various times and scan-angles. If the separation is small the blending of the sources corresponds to a simple shift in the inferred position. As we approach large separations ($\rho\gtrsim 1$) the asymmetry of the effective width $\gamma$ introduces extra angular dependence which cannot be captured by a simple offset and leads to excess noise on astrometric fits. Depending on the pattern of scan angles with time this noise may translate into a bias in proper motion or parallax (for further discussion see El-Badry24b). Blends between unrelated background and foreground stars will show the same behaviour, provided that their astrometric motion (parallax and proper motion) are similar or negligible.
@@
Thus a population of long period semi-resolved binaries also contributes to the known population of binaries with significant internal astrometric motion (periods of order of the timespan of the observations, of order years for Gaia Penoyre22). With epoch astrometry data (as will be available in the next Gaia data release) these two cases can be differentiated, but currently they cannot easily be separated, even though they represent markedly different types of binary.
@@ Appendix A Maximal photocentre offsets
Figure 8: Dark red: maximum possible displacement of the primary peak, $\chi_{max}=x_{max}/\gamma$, as a function of $l$. Light red: the overestimated value that would be derived from assuming $\chi=\rho l/(1+l)$. Grey: the limiting value at low $l$ of $le^{-\frac{1}{2}}$.
@@ Appendix B Approximate form for low $\rho$
$x_{m}\approx x_{0}\left(1+\frac{x_{0}^{2}}{l\gamma^{2}}\right)$ (58)
@@ B.1 Offset and error
$k=x_{0}+\frac{x_{0}^{3}}{l}\frac{\left\langle\gamma^{-4}\right\rangle}{\left\langle\gamma^{-2}\right\rangle}+\mathcal{O}(\rho^{4})$ (59)
@@
$k=x_{0}+\frac{x_{0}^{3}}{4l\alpha^{2}}\frac{3+2\eta+3\eta^{2}}{1+\eta}+\mathcal{O}(\rho^{4})$ (64)
@@
For a circular PSF, $\eta=1$, $k$ goes to its maximum value of $x_{m}$ (as given by equation 58) and $D=0$ as here the image has no angular dependence. Reducing $\eta$, starting from a value of 1, $k$ decreases and $D$ increases, with both reaching their limiting value at $\eta=\frac{2}{\sqrt{3}}-1\approx 0.1547$ where $D\approx 0.03775\frac{x_{0}^{6}}{l^{2}\alpha^{6}}$. The existence of a $\eta$ which maximises $D$ suggests the possibility of an optimal instrument design for detecting the presence of sub-resolution binaries in images (or alternatively a worst-case design if one wanted to avoid the excess noise that such sources cause). As one final interesting serendipity, the Gaia mission design, with $\eta\approx\frac{1}{3}$ gives almost identical $k$ and $D$ as the $\eta=0$ case (though this ceases to be true outside of the $\rho\ll 1$ case).

### v2

@@ 1 Introduction
This applies to pairs of sources that are truly close in physical space (e.g. perhaps a gravitationally bound binary system) or that are at very different distances from the observer and each other but appear close in projection. It is also eminently possible for more than two sources to be blended - for example whole clusters of stars appear as a bright point source in extragalactic observations (Harris, 1991). In fact we may expect many blended sources to be contributing to an observed image (for example planets and asteroids around a star, or many more distant background stars) though if these systems contribute a negligible fraction of the observed light, their influence is small.
@@
Binary stars (and higher multiples) are ubiquitous in most astrophysical environments. Only very wide systems, or those very close to the observer, are resolvable - however multiplicity affects many facets of their behaviour and evolution. To some they are objects of interest in their own right, to others a nuisance to be removed or compensated for when making observations. Only a subset can be identified, and only a subset of those can have their properties measured. The number of known multiples is rapidly increasing, both by resolving wide systems (for example Tokovinin 2024) or, for unresolved systems, by identifying variable or excess motion (spectroscopically or astrometrically, see El-Badry 2024 for a recent review) or light (e.g. Prša and Zwitter 2005 or Wallace 2024).
The Gaia survey has been crucial to the improvement in our capacity to detect stellar multiples, as well as greatly expanding our knowledge of the position and motion of stars in the Milky Way (Gaia Collaboration et al., 2016, 2023). It is, at least in part, an astrometric survey - scanning repeatedly over the sky and making many measurements of the positions of sources, from which their motion can be inferred (Lindegren et al., 2021). In Penoyre et al. (2020) we mapped out how the orbits of unresolved binaries contribute to this motion, due to the offset between their photocentre (or center-of-light) and their center-of-mass. In appendix A of that work, we derived the simple first order photocentre offset distance, without including an important nuance to Gaia’s functionality: that its mirror is rectangular, measuring 1.45 m by 0.5 m (Gaia Collaboration et al., 2016), and hence its subsequent PSF is extended along the direction corresponding to the short axis of the mirror.
@@ 2 Brightness profiles
Throughout this work we will work with Gaussian point-spread functions (PSF). These are vastly simplified from real diffraction patterns. Idealised forms of elongated PSFs include convolutions of two $\mathrm{sinc}^{2}$ functions for a rectangular aperture, or an elongated form of the well known Airy disk for an elliptical aperture (Airy, 1835; Raman, 1919).
@@
True PSFs will be further altered by any non-ideal aberrations to the mirror, optical path, or detector. For example Fabricius et al. (2016) show a typical Gaia PSF (figure 7) which as expected resembles a smoothed $\mathrm{sinc}^{2}$ function, but displays significant asymmetry and strong diagonal spikes. Real observations are then further altered by pixelation, losing more information as the incident flux is binned over a finite number of pixels. Very bright sources can oversaturate these pixels, spilling charge and thus apparent brightness into neighbouring pixels, and further deforming the PSF.
@@ 2.1 Imaging close sources
For derivations in this paper we’ll work entirely in units of $\alpha$ and $\beta$ (and hence we shall not need numerical values specific to a given instrument). However, as Gaia is the most currently relevant point of comparison we note that its full-width-half-maximum is approximately 0.1 arcseconds (Fabricius et al., 2016) in the along scan direction, and hence $\alpha\sim 0.04$ arcseconds. Given that the ratio of side lengths of the mirror is 1.45:0.5 we expect $\beta\sim 3\alpha$.
@@
Real observations may use a more complex PSF fitting technique (e.g. Rowell et al. 2021; Harrison et al. 2023), but the above assumptions provide a reasonable approximation that is well defined mathematically.
@@ 2.2 Maxima and minimum
Equations 14 and 15 are the equivalent of equation 64 in Lindegren (2022). However, their calculation only includes the along-scan PSF width, $\alpha$ ($u$ in their notation), implicitly assuming $\beta\gg\alpha$ (and hence $\gamma\rightarrow\alpha/|\mathrm{c}_{\phi}|$). This significantly overestimates the effective width as $|\mathrm{c}_{\phi}|\rightarrow 0$.
@@ 2.2.1 Approximate forms
where $x_{0}$ is the much simplified zeroth order equation for the photocentre reported in Penoyre et al. (2020).
@@
$x_{m}=x_{0}\left(1-\frac{x_{0}^{2}}{l\gamma^{2}}\right)+\mathcal{O}\left(\frac{r^{4}}{\gamma^{4}}\right).$ (18)
@@ 3.1 Numerical solutions
We seek the solution(s), $x_{m}$, of equation 13. This can be expressed as root(s) of $g(\chi_{m})=0$ which can be reliably found by any root-solver which operates over a range $a<\chi<b$ such that $g(a)$ and $g(b)$ have opposite sign and there is a single root in this range (for example the van Wijngaarden-Deker-Brent method, Brent 1973 - implemented as brentq in scipy.optimize).
@@
The numerically derived position of each root is shown for a variety of $l$ and $\rho$ in figure 4. For small separations ($\rho<1$) the primary maximum is located close to the first order solution as found in Penoyre et al. (2020): $x_{m}\sim x_{0}$. As $\rho$ increases the relative offset decreases, fastest for smaller $l$ - thus the first order solution is a good approximation for small $\rho$ and an overestimate for $\rho\gtrsim$ 1. In appendix A we derive the conditions that maximise $x_{1}$, as is also shown in the figure. At low $l$ the photocenter offset is maximised at $\rho=1$, and as $l$ increases this goes to $\rho=2$.
@@ 4 Repeated observations of a fixed pair of sources
We also show the across-scan position that one would find if $\beta\rightarrow\infty$ - as is implicitly assumed by the calculation in Lindegren (2022). This would lead to greater predicted offsets, especially at large separations, as any source observed at an angle such that $|\mathrm{c}_{\phi}|\lesssim\alpha/r$ would appear to be completely blended.
@@ 4.1 Binary offset and error
The precision to which the position of a source can be measured depends on the width of the PSF and how well sampled it is. For $N$ samples (i.e. measured photons) of a distribution with PSF width $\sigma$ we might expect a positional uncertainty of $\sim N^{-\frac{1}{2}}\sigma$. Other observational constraints, such as saturation of CCD pixels, can limit the possible precision. Lindegren et al. (2021) (specifically appendix A) detail the current along-scan astrometric precision of Gaia, showing that it is photon noise dominated for systems with apparent magnitude $m_{G}\gtrsim 14$, and for brighter sources is approximately constant, with $\sigma_{al}\sim 0.2$ milliarcseconds.
@@ 4.2 Observational expectations
Assuming a single source (as the main Gaia source catalogue does) this motion is a combination of proper motion and parallax. If the scans are frequent and randomly oriented we would not expect any consistent behaviour that could mimic astrometric motion, thus we will only consider the position offset in detail in this work. However if there is some aliasing or coherent time evolution of scan-angles and times the varying positional offset may mimic motion, as seen in Holl et al. (2023).
If there is excess unmodelled astrometric motion, for example from binarity, there will be a poor astrometric fit. Due to the ubiquity of binaries and higher multiples (accounting for around half of all systems Raghavan et al. 2010) persistently poor fits can be used to identify probable multiple systems (see e.g. Penoyre et al. 2020; Belokurov et al. 2020; Penoyre et al. 2022b).
@@
In Gaia DR3 a system can be distinguished as a significantly bad fit (whether due to internal motion or other contaminants) if the excess $RUWE$ is $\gtrsim 0.25$ and thus (Penoyre et al., 2022a; Castro-Ginard et al., 2024).
@@ 4.2.1 Observable offset and error
We see the largest deviations from good astrometric fits, $\sqrt{1+D}-1$, at high light ratios and in the partially resolved region. The excess noise can be very large, well above the threshold of 0.25 appropriate for current Gaia data. Even excluding the semi-resolved region excesses of up to 50 are possible. This is a relatively hard predicted upper limit, and agrees well with the maximum RUWE seen in sources separated by around 1 arcseond (Belokurov et al. 2020, figure 5).
@@ 4.2.2 Observational complications
Systems in the offset and error regime are the most amenable to detection with other methods, more observations or in later surveys. The large light ratios make them easier to discern from single sources, either from elongated photometry, or two discernable profiles in their spectra. Gaia includes flags based on the Image Parameter Determination (IPD) pipeline that record deviations from expected point-source brightness profiles which can be used to flag some sources in this regime (see for example appendix B of Penoyre et al. 2022b). The pipeline for Gaia DR4 will identify and fit secondary peaks where possible, though the regime of interest here is at the edge of, or below, the threshold of resolvability and thus this will mitigate the effects mostly of the partially resolved systems.
@@ 5 Conclusions
We start, in section 2, with an elongated Gaussian PSF, as is appropriate for surveys such as the Gaia mission, with a rectangular focussing mirror. We show that this more general case can be reduced to a 1D problem, and thus converges with the circular PSF case when normalised by a scan-angle dependant effective PSF width, $\gamma(\phi)$ (equation 16). Thus any general case can be expressed in terms of $\rho=\frac{r}{\gamma}$, where $r$ is the on-sky separation of the sources. This introduces a further dependence on the relative orientation of the sources and our instrument, and shows that when this orientation changes over multiple measurements the variation in observed photocentre can provide extra information about the relative separation and light ratio of the sources. Such variations have indeed been observed in Gaia already, as detailed in Holl et al. (2023).
@@
When two blended sources maintain a fixed orientation analytic expressions for their position offset and induced error can be derived, as we do in section 4. This is a particular case of interest for long period (and hence effectively stationary) binaries observed at various times and scan-angles. If the separation is small the blending of the sources corresponds to a simple shift in the inferred position. As we approach large separations ($\rho\gtrsim 1$) the asymmetry of the effective width $\gamma$ introduces extra angular dependence which cannot be captured by a simple offset and leads to excess noise on astrometric fits. Depending on the pattern of scan angles with time this noise may translate into a bias in proper motion or parallax (for further discussion see El-Badry et al. 2024). Blends between unrelated background and foreground stars will show the same behaviour, provided that their astrometric motion (parallax and proper motion) are similar or negligible.
@@
Thus a population of long period semi-resolved binaries also contributes to the known population of binaries with significant internal astrometric motion (periods of order of the timespan of the observations, of order years for Gaia Penoyre et al. 2022a). With epoch astrometry data (as will be available in the next Gaia data release) these two cases can be differentiated, but currently they cannot easily be separated, even though they represent markedly different types of binary.
@@ Appendix A Maximal photocentre offsets
Figure 8: Solid line: maximum possible displacement of the primary peak, $\chi_{max}=x_{max}/\gamma$, as a function of $l$. Dashed line: the overestimated value that would be derived from assuming $\chi=\rho l/(1+l)$. Dotted line: the limiting value at low $l$ of $le^{-\frac{1}{2}}$.
@@ Appendix B Approximate form for low $\rho$
$x_{m}\approx x_{0}\left(1-\frac{x_{0}^{2}}{l\gamma^{2}}\right)$ (58)
@@ B.1 Offset and error
$k=x_{0}-\frac{x_{0}^{3}}{l}\frac{\left\langle\gamma^{-4}\right\rangle}{\left\langle\gamma^{-2}\right\rangle}+\mathcal{O}(\rho^{4})$ (59)
@@
$k=x_{0}-\frac{x_{0}^{3}}{4l\alpha^{2}}\frac{3+2\eta+3\eta^{2}}{1+\eta}+\mathcal{O}(\rho^{4})$ (64)
@@
For a circular PSF, $\eta=1$, $k$ goes to its minimum value of $x_{m}$ (as given by equation 58) and $D=0$ as here the image has no angular dependence. Reducing $\eta$, starting from a value of 1, $k$ and $D$ increase, with both reaching their limiting value at $\eta=\frac{2}{\sqrt{3}}-1\approx 0.1547$ where $D\approx 0.03775\frac{x_{0}^{6}}{l^{2}\alpha^{6}}$. The existence of a $\eta$ which maximises $D$ suggests the possibility of an optimal instrument design for detecting the presence of sub-resolution binaries in images (or alternatively a worst-case design if one wanted to avoid the excess noise that such sources cause). As one final interesting serendipity, the Gaia mission design, with $\eta\approx\frac{1}{3}$ gives almost identical $k$ and $D$ as the $\eta=0$ case (though this ceases to be true outside of the $\rho\ll 1$ case).